I need verification with this equation please.
√(n-4) + √(n+4)= 2√(n-1)
I keep getting {0,1} which are not valid
but the answer in back of my book says the answer is {5}, I don't get it.
please help thank you

Answer by lwsshak3(11628)

√(n-4) + √(n+4)= 2√(n-1)
square both sides
(n-4)+2√(n-4)*√(n+4)+(n+4)=4(n-1)
n-4+2√(n-4)(n+4)+n+4=4n-4
2√(n-4)(n+4)+2n=4n-4
2√(n^2-16)=2n-4
square both sides again
4(n^2-16)=4n^2-16n+16
4n^2-64=4n^2-16n+16
16n=80
n=5
ans:{5}